Emma has half of her investments in a stock paying 10% dividends and the other half in a stock paying 15% interest. If her total annual interest is $510, how much does she have invested in each and total?
Answer by mananth(16946)

let the investment in each be x
total investment = 2x
10%x+15%x=510
0.1x+0.15x=510
0.25x=510
x= 510/0.25
x= 2040 investment in each
total = 2040 *2= 4080